# Diversity-Oriented Data Augmentation with Large Language Models

### v1

@@ Abstract
Data augmentation is an essential technique in natural language processing (NLP) for enriching training datasets by generating diverse samples. This process is crucial for improving the robustness and generalization capabilities of NLP models. However, a significant challenge remains: Insufficient Attention to Sample Distribution Diversity. Most existing methods focus on increasing the sample numbers while neglecting the sample distribution diversity, which can lead to model overfitting. In response, we explore data augmentation’s impact on dataset diversity and propose a Diversity-oriented data Augmentation framework (DoAug). Specifically, we utilize a diversity-oriented fine-tuning approach to train an LLM as a diverse paraphraser, which is capable of augmenting textual datasets by generating diversified paraphrases. Then, we apply the LLM paraphraser to a selected coreset of highly informative samples and integrate the paraphrases with the original data to create a more diverse augmented dataset. Finally, we conduct extensive experiments on 12 real-world textual datasets. The results show that our fine-tuned LLM augmenter improves diversity while preserving label consistency, thereby enhancing the robustness and performance of downstream tasks. Specifically, it achieves an average performance gain of $10.52\%$, surpassing the runner-up baseline with more than three percentage points.
@@
Zaitian Wang1,2, Jinghan Zhang3, Xinhao Zhang3, Kunpeng Liu3, Pengfei Wang1,2, Yuanchun Zhou1,2 1Computer Network Information Center, CAS, 2University of Chinese Academy of Sciences, 3Portland State University wangzaitian23@mails.ucas.ac.cn
@@ 1 Introduction
AI methods have demonstrated immense capabilities, often surpassing human abilities and traditional techniques across various natural language processing (NLP) tasks. This success largely hinges on the availability of high-quality datasets, which enable AI models to uncover intrinsic patterns and drive their effectiveness in real-world applications. However, training on inferior datasets can significantly degrade model performance, particularly when applied to test data or real-world scenarios Wang et al. (2024b). As AI technology advances, especially with large language models (LLMs), the demand for high-quality datasets has become more pronounced. To effectively train NLP models, a high-quality dataset should be (1) Large: a sufficient number of samples is crucial to reflect the diversity and complexity of human language. Large datasets help prevent overfitting, ensuring the trained AI model generalizes well to unseen data. With more data points, the model can learn various patterns and relationships, which enhances its robustness and reliability; (2) Coherent: the mapping between data and labels must be accurate and consistent. Coherent datasets ensure that each data point is correctly labeled, providing the model with reliable information for learning. Incoherent datasets, with mislabeled or inconsistent data, can confuse the model and degrade its performance. Consistency in labeling also aids in the reproducibility of results and the interpretability of the model’s predictions; (3) Diverse: a diverse dataset ensures NLP models learn a broad spectrum of linguistic patterns, enhancing robustness across real-world conditions. This includes variations such as dialects, tones, formality, or domain-specific terms. Exposure to such diversity helps models generalize better, avoiding over-reliance on narrow language subsets. Additionally, it improves adaptability to unexpected inputs while reducing biases tied to certain language styles.
@@
Data augmentation is an efficient technique for increasing the number of training samples by modifying existing dataset samples Wang et al. (2024b). It allows for the rapid generation of large-scale datasets without the need for additional data collection and has been successfully applied in various domains, including textual data. Data augmentation for textual data often changes the wording or reshapes the structure of a sentence. Many early and simple methods achieve this by randomly perturbing the textual samples at the word level Wei and Zou (2019); Karimi et al. (2021). While certain operations prove effective, some introduce noise that compromises label integrity (e.g. deleting a “not”) or generates almost redundant samples, ultimately failing to improve dataset quality or promote diversity. Recent advances in LLMs have demonstrated unprecedented power in text understanding and generation Radford et al. (2018). The capacity of these state-of-the-art (SOTA) generative language models establishes a new and promising paradigm of textual data augmentation Anaby-Tavor et al. (2020). Generative models enable large-scale acquisition of textual data while preserving the coherence of augmented datasets by generating texts with similar meanings to the original sentences Dai et al. (2025). However, most existing generative methods focus primarily on enlarging dataset size, with limited consideration of how the augmentation process affects dataset diversity. Adequate attention to maintaining and enhancing dataset diversity is vital for developing AI-ready, high-quality datasets, making it a crucial challenge in textual data augmentation.
@@
Along this line, we propose a Diversity-oriented data Augmentation approach (DoAug) using an LLM to paraphrase sentences. The LLM is first fine-tuned on a paraphrase dataset and taught to rewrite sentences. By instructing the LLM to function as a paraphraser, we can use it to alter sentence expressions while preserving their essential meaning. In this way, we ensure the affinity between the original and augmented samples, minimizing the influence of data augmentation on dataset coherence. To enhance the dataset diversity through data augmentation, we further proposed a diversity-oriented fine-tuning method. We construct a preference dataset that chooses the more diverse paraphrases while rejecting repetitive ones. Then the LLM paraphraser is fine-tuned on the preference dataset with the DPO algorithm Rafailov et al. (2024) to encourage greater diversity in its generated samples. We also adopt a coreset selection method to focus on only the most important samples from the dataset to reduce the computational overhead and costs of running LLMs. Finally, we conduct extensive experiments on 12 textual benchmark datasets to verify the effectiveness of DoAug. Experimental results show that our proposed method can remarkably enhance the diversity of the augmented dataset on 5 measurements while maintaining high affinity compared with the original dataset. We also investigate the implications of this increased diversity on model performance in downstream tasks and observe significant improvements in model performance when trained on the augmented datasets.
@@
(3) We introduce a diversity-oriented fine-tuning method that trains the LLM augmenter on a preference dataset with the DPO algorithm the boost generation diversity of the LLM;
@@ 2.1 Textual Data Augmentation
Textual data augmentation revolves around perturbing the wording and syntax of existing sentences to create more modified samples. Some early and simple methods propose to randomly replace, remove, insert, and swap characters or words at certain ratios in a sentence Belinkov and Bisk (2018); Wei and Zou (2019). Some more sophisticated methods modify sentences by using alternative syntax Min et al. (2020). Language models can in turn act as effective tools for textual data augmentation. For example, Back-translation Sennrich (2015) first translates sentences from the source language (e.g. English) to an intermediary (e.g. Chinese) and then translates the intermediary sentence back to the source language. Substitute Word using BERT Kumar et al. (2020) masks certain words in the original sentences and uses a BERT model to predict masked words. They utilize the subtle differences made by the translator or the unmasking process and perturb the wording or syntax while keeping the meanings untouched. The recent emergence of LLM has given birth to a series of new approaches Anaby-Tavor et al. (2020); Cai et al. (2023); Ding et al. (2024); Wang et al. (2024a). AugGPT Dai et al. (2025), for example, prompts the state-of-the-art ChatGPT model to rewrite sentences in the dataset and preserves dataset coherence after data augmentation. Self-LLMDA Li et al. (2024) automatically generates and selects the most suitable instruction to prompt the LLM to generate augmented samples. However, these aforementioned methods neglect the impact on dataset diversity, failing to ensure the diversity trait of producing high-quality datasets. The effect of LLM augmentation diversity is discussed in Cegin et al. (2023, 2024), where three types of prompt-based diversity incentives are proposed.
@@ 3.1 Problem Formulation
$\vspace{-3mm}\theta^{\ast}=\mathop{\arg\max}_{\theta}\mathbbm{E}\Delta D(\mathcal{S};\theta)$ (1)
@@
$\vspace{-3mm}\Delta P(\mathcal{S};\theta^{\ast})=\mathop{\max}_{\theta}\Delta P(\mathcal{S};\theta)$ (2)
@@ 3.2 Framework Overview
DoAug trains and employs an LLM capable of generating diverse paraphrases for data augmentation to enlarge the size of textual datasets, maintain coherence, and enhance diversity. As Figure 1 shows, the framework is organized as follows:
@@ 3.3.2 LLM Generation Diversity Enhancement with DPO
To align LLM generations to human preferences, recent LLMs adopt RLHF in their training process, which involves the PPO algorithm and a reward model Ouyang et al. (2022); Zhang et al. (2024). However, fitting a reward model brings extra computation costs, and the gap between its prediction and actual human preference also poses threats to the effect of PPO. As an alternative, the DPO algorithm directly optimizes the LLMs’ generation policy without training an additional reward model Rafailov et al. (2024).
@@
$\vspace{-3mm}\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{DPO}}}\bigg{[}\log\sigma\bigg{(}\\ \beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\left.\left.-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (6)
@@ 3.3.3 Diversity-oriented Sampling
8:  Rank samples in $\mathcal{S}^{\prime}$ according to the score key $s$
@@
16:  $\mathcal{\tilde{S}}=\mathcal{\tilde{S}}\cup\mathcal{S}_{\text{select}}$
@@ 4.1 Experiment Settings
Evaluation Criterion. We evaluate diversity and affinity for data distribution while measuring performance on downstream tasks for effectiveness.
Diversity. To comprehensively evaluate the effect of the proposed model DoAug on dataset diversity, we adopt several measures to evaluate the augmented dataset’s diversity:
$\bullet$ Distance: assesses the average distance between samples as follows: ${Distance(\mathcal{S})=\frac{1}{|\mathcal{S}|}\sum_{x_{i},x_{j}\in\mathcal{S}}\sqrt{(e_{x_{i}}-e_{x_{j}})^{2}}}$, where $e_{x}=\mathcal{E}(x)$ is the embedding of sample $x$ in the embedding space $\mathcal{E}$, and a larger distance indicates greater diversity.
[…]
@@
Datasets. We conduct extensive experiments on 12 NLP datasets to verify the effectiveness of DoAug. As specified in Appendix A, our selection of datasets covers a wide range of text classification tasks and includes datasets with different characteristics. Following the settings in Yoo et al. (2021), we sample a subset (1.2K samples) from the full dataset to unify the evaluation settings, enable a fair comparison between methods, and simulate a low-resource condition where data augmentation is of significant necessity.
Baseline Methods. We compare DoAug with eleven representative data augmentation methods. (1) OCR and (2) Keyboard perform common OCR or typing errors at the character level. (3) EDA randomly inserts, deletes, replaces, or swaps words in the sentences to create more samples. (4) Back-translation (BT) involves translating the source sentences to an intermediary language. (5) Unmask randomly replaces words with [MASK] and predicts the masked words with the BERT model. (6) AugGPT directly prompts ChatGPT for paraphrases without further fine-tuning Dai et al. (2025). (7) Grammar and (8) Spelling are two exemplar methods selected by Self-LLMDA Li et al. (2024), which prompt the LLM to simulate common grammatical variation or spelling errors made by humans. (9) Chain, (10) Hint, and (11) Taboo generate paraphrases with three different diversity incentives Cegin et al. (2024).
@@
Figure 2: Diversity, affinity, and performance achieved by DoAug and baseline methods. Results are averaged on 12 datasets and the diversity rankings are further averaged on 5 metrics in this diagram. Smaller number for the rankings indicates better results.

### v2

@@ Abstract
Data augmentation is an essential technique in natural language processing (NLP) for enriching training datasets by generating diverse samples. This process is crucial for improving the robustness and generalization capabilities of NLP models. However, a significant challenge remains: Insufficient Attention to Sample Distribution Diversity. Most existing methods focus on increasing the sample numbers while neglecting the sample distribution diversity, which can lead to model overfitting. In response, we explore data augmentation’s impact on dataset diversity and propose a Diversity-oriented data Augmentation framework (DoAug). Specifically, we utilize a diversity-oriented fine-tuning approach to train a large language model (LLM) as a diverse paraphraser, which is capable of augmenting textual datasets by generating diversified paraphrases. Then, we apply the LLM paraphraser to a selected coreset of highly informative samples and integrate the paraphrases with the original data to create a more diverse augmented dataset. Finally, we conduct extensive experiments on 12 real-world textual datasets. The results show that our fine-tuned LLM augmenter improves diversity while preserving label consistency, thereby enhancing the robustness and performance of downstream tasks. Specifically, it achieves an average performance gain of $10.52\%$, surpassing the runner-up baseline with more than three percentage points.
@@
Zaitian Wang1,2, Jinghan Zhang3, Xinhao Zhang3, Kunpeng Liu3, Pengfei Wang1,2111Corresponding author, Yuanchun Zhou1,2 1Computer Network Information Center, CAS, 2University of Chinese Academy of Sciences, 3Portland State University wangzaitian23@mails.ucas.ac.cn
@@ 1 Introduction
Figure 1: Conceptual comparison of DoAug (right) generating coherent and diverse samples against baselines (left) generating noisy or repetitive samples.
@@
AI methods have demonstrated immense capabilities, often surpassing human abilities and traditional techniques across various natural language processing (NLP) tasks. This success largely hinges on the availability of high-quality datasets, which enable AI models to uncover intrinsic patterns and drive their effectiveness in real-world applications Xu et al. (2025); Zhang et al. (2025). However, training on inferior datasets can significantly degrade model performance, particularly when applied to test data or real-world scenarios Wang et al. (2024b). As AI technology advances, especially with large language models (LLMs), the demand for high-quality datasets has become more pronounced Wang et al. (2025). To effectively train NLP models, a high-quality dataset should be (1) Large: a sufficient number of samples is crucial to reflect the diversity and complexity of human language. Large datasets help prevent overfitting, ensuring the trained AI model generalizes well to unseen data. With more data points, the model can learn various patterns and relationships, which enhances its robustness and reliability; (2) Coherent: the mapping between data and labels must be accurate and consistent. Coherent datasets ensure that each data point is correctly labeled, providing the model with reliable information for learning. Incoherent datasets, with mislabeled or inconsistent data, can confuse the model and degrade its performance. Consistency in labeling also aids in the reproducibility of task results and the interpretability of the model’s predictions; (3) Diverse: a diverse dataset ensures NLP models learn a broad spectrum of linguistic patterns, enhancing robustness across real-world conditions. This includes variations such as dialects, tones, formality, or domain-specific terms. Exposure to such diversity helps models generalize better, avoiding over-reliance on narrow language subsets. Additionally, it improves adaptability to unexpected inputs while reducing biases tied to certain language styles.
@@
Data augmentation is an efficient technique for increasing the number of training samples by modifying existing dataset samples Wang et al. (2024b). It allows for the rapid generation of large-scale datasets without the need for additional data collection and has been successfully applied in various domains, including textual data. Data augmentation for textual data often changes the wording or reshapes the structure of a sentence. Many early and simple methods achieve this by randomly perturbing the textual samples at the word level Wei and Zou (2019); Karimi et al. (2021). While certain operations prove effective, some are prone to introducing noise that compromises label integrity (e.g. deleting a “not”) or generating redundant samples, ultimately failing to improve dataset quality or promote diversity (Figure 1 left). Recent advances in LLMs have demonstrated unprecedented power in text understanding and generation Radford et al. (2018); Chen et al. (2024). The capacity of these state-of-the-art (SOTA) generative language models establishes a new and promising paradigm of textual data augmentation Anaby-Tavor et al. (2020). Generative models Brown et al. (2020); Ning et al. (2024); Dubey et al. (2024) enable large-scale acquisition of textual data while preserving the coherence of augmented datasets by generating texts with similar meanings to the original sentences Dai et al. (2025). However, most existing generative methods focus primarily on enlarging dataset size, with limited consideration of how the augmentation process affects diversity. Adequate attention to maintaining and enhancing dataset diversity is vital for developing AI-ready and high-quality datasets, making it a crucial challenge in designing textual data augmentation methods.
@@
Along this line, we propose a Diversity-oriented data Augmentation approach (DoAug) using an LLM to paraphrase sentences (Figure 1 right). The LLM is first fine-tuned on a paraphrase dataset and taught to rewrite sentences. By instructing the LLM to function as a paraphraser, we can use it to alter sentence expressions while preserving their essential meaning. In this way, we ensure the affinity between the original and augmented samples, minimizing the influence of data augmentation on dataset coherence. To enhance the dataset diversity through data augmentation, we further proposed a diversity-oriented fine-tuning method. We construct a preference dataset that chooses the more diverse paraphrases while rejecting repetitive ones. Then the LLM paraphraser is fine-tuned on the preference dataset with the DPO algorithm Rafailov et al. (2024) to encourage greater generation diversity. We also adopt a coreset selection method to focus on only the most important samples from the dataset to reduce the computational overhead and costs of running LLMs. Finally, we conduct extensive experiments on 12 textual benchmark datasets to verify the effectiveness of DoAug. Experimental results show that our proposed method can remarkably enhance the diversity of the augmented dataset on 6 measurements while maintaining high affinity compared with the original dataset. We also investigate the implications of this increased diversity on model performance in downstream tasks and observe significant improvements in model performance when trained on the augmented datasets.
@@
(3) We introduce a diversity-oriented fine-tuning method that trains the LLM augmenter on a preference dataset with the DPO algorithm to boost the generation diversity of the LLM;
@@ 2.1 Textual Data Augmentation
Textual data augmentation revolves around perturbing the wording and syntax of existing sentences to create more modified samples. Some early and simple methods propose to randomly replace, remove, insert, and swap characters or words at certain ratios in a sentence Belinkov and Bisk (2018); Wei and Zou (2019). Some more sophisticated methods modify sentences by using alternative syntax Min et al. (2020). Language models can in turn act as effective tools for textual data augmentation. For example, Back-translation Sennrich et al. (2016) first translates sentences from the source language (e.g. English) to an intermediary (e.g. Chinese) and then translates the intermediary sentence back to the source language. Substitute Word using BERT Kumar et al. (2020) masks certain words in the original sentences and uses a BERT model to predict masked words. They utilize the subtle differences made by the translator or the unmasking process and perturb the wording or syntax while keeping the meanings untouched. The recent emergence of LLM has given birth to a series of new approaches Anaby-Tavor et al. (2020); Cai et al. (2023); Ding et al. (2024); Wang et al. (2024a). AugGPT Dai et al. (2025), for example, prompts the state-of-the-art ChatGPT model to rewrite sentences in the dataset and preserves dataset coherence after data augmentation. Self-LLMDA Li et al. (2024) automatically generates and selects the most suitable instruction to prompt the LLM to generate augmented samples. However, these aforementioned methods neglect the impact on dataset diversity, failing to ensure the diversity trait of producing high-quality datasets. The effect of LLM augmentation diversity is discussed in Cegin et al. (2023, 2024), where three types of prompt-based diversity incentives are proposed.
@@ 3.1 Problem Formulation
$\theta^{\ast}=\mathop{\arg\max}_{\theta}\mathbbm{E}\Delta D(\mathcal{S};\theta)$ (1)
@@
$\Delta P(\mathcal{S};\theta^{\ast})=\mathop{\max}_{\theta}\Delta P(\mathcal{S};\theta)$ (2)
@@ 3.2 Framework Overview
DoAug trains and employs an LLM capable of generating diverse paraphrases for data augmentation to enlarge the size of textual datasets, maintain coherence, and enhance diversity [^1] Code is available at: https://github.com/CNICDS/DoAug. As Figure 2 shows, the framework is organized as follows:
@@ 3.3.2 LLM Generation Diversity Enhancement with DPO
To align LLM generations to human preferences, recent LLMs adopt RLHF in their training process, which involves the PPO algorithm and a reward model Ouyang et al. (2022); Zhang et al. (2024b, a). However, fitting a reward model brings extra computation costs, and the gap between its prediction and actual human preference also poses threats to the effect of PPO. As an alternative, the DPO algorithm directly optimizes the LLMs’ generation policy without training an additional reward model Rafailov et al. (2024).
@@
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{DPO}}}\bigg{[}\log\sigma\bigg{(}\\ \beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\left.\left.-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (6)
@@ 3.3.3 Diversity-oriented Sampling
8:  Rank samples in $\mathcal{S}^{\prime}$ according to the score $s$
@@
16:  $\mathcal{\tilde{S}}=\mathcal{\tilde{S}}\cup\mathcal{S}_{\text{retain}}$
@@ 4.1 Experiment Settings
4.1.1 Evaluation Criterion
We evaluate diversity and affinity for data distribution while measuring performance on downstream tasks for effectiveness. For all these measurements, the higher score indicates better results.
Diversity. To comprehensively evaluate DoAug’s effect on dataset diversity, we adopt several measurements to assess the augmented dataset’s diversity from both latent and lexical aspects:
$\bullet$ Distance and Dispersion assess datasets’ latent diversity at the sample level by calculating pair-wise Euclidean distance and Cosine similarity on the embedding space.
$\bullet$ Isocontour Radius and Homogeneity assess datasets’ latent diversity at the dataset level by considering the coverage and uniformity of all sample embeddings.
$\bullet$ Vocabulary Size and Unique 3-grams assess datasets’ lexical diversity by counting how many different words are used throughout the datasets.
Affinity. The affinity score reflects the coherence of an augmented dataset and is embodied by embedding deviation.
@@ 4.1.1 Evaluation Criterion
4.1.2 Datasets
We conduct extensive experiments on 12 NLP datasets to verify the effectiveness of DoAug. Our selection of datasets covers a wide range of text classification tasks, including entailment annotation (ANLI, MNLI, and RTE), sentiment analysis (MPQA, SST-2, and Yelp), chemical-protein relationship (ChemProt), acceptability judgment (CoLA), semantically equivalence (MRPC), sentence role (RCT), subjectivity analysis (SUBJ), and Disease judgment (Symptoms) Pang and Lee (2004); Wiebe et al. (2005); Zhang et al. (2015); Kringelum et al. (2016); Dernoncourt and Lee (2017); Wang et al. (2019); Nie et al. (2020); Dai et al. (2025). More details of these datasets are specified in Appendix B. Following the settings in Yoo et al. (2021), we sample a subset (1.2K samples) from the full dataset to unify the evaluation settings, enable a fair comparison between methods, and simulate a low-resource condition where data augmentation is of significant necessity.
4.1.3 Baseline Methods
We compare DoAug with twelve representative data augmentation methods. (1) OCR and (2) Keyboard perform common OCR or typing errors at the character level Li et al. (2024). (3) EDA randomly inserts, deletes, replaces, or swaps words in the sentences Wei and Zou (2019). (4) AEDA randomly inserts punctuations in the sentences Karimi et al. (2021). (5) Back-translation (BT) involves translating the source sentences to an intermediary language Sennrich et al. (2016). (6) Unmask randomly replaces words with [MASK] and predicts the masked words with the BERT model Kumar et al. (2020). (7) AugGPT directly prompts ChatGPT (replaced with Llama3.1-8B-Instruct to save cost) for paraphrases without further fine-tuning Dai et al. (2025). (8) Grammar and (9) Spelling are two exemplar methods selected by Self-LLMDA, which prompt the LLM to simulate common grammatical variation or spelling errors made by humans Li et al. (2024). (10) Chain, (11) Hint, and (12) Taboo generate paraphrases with three different diversity incentives Cegin et al. (2024). Augmentation examples of these methods are presented and discussed in Appendix E Table 6.
4.1.4 Key Implementation Information
We use Llama-3.2-1B-Instruct with BF16 quantization as the LLM paraphraser. The LLM’s training dataset $\mathcal{D}_{\text{SFT}}$ and $\mathcal{D}_{\text{DPO}}$ contain 100,000 sentence pairs and 50,000 preference pairs, respectively. In the coreset selection step, the coreset ratio $r_{\text{augment}}:r_{\text{retain}}:r_{\text{prune}}$ is $1:1:1$. We explain how we derive this ratio in Appendix I.2. For each original sentence, we generate $K=5$ paraphrases and sample the most diversified output. A more detailed implementation specification is given in Appendix C.
@@ 4.1.4 Key Implementation Information
Distance Dispersion Radius Homogeneity Vocabulary 3-grams Average
Original 0.00 0.00 0.78 0.74 0.00 0.00 0.25
OCR 0.01 0.05 0.62 0.83 0.05 0.11 0.28
Keyboard 0.00 0.05 0.55 0.83 0.08 0.17 0.28
EDA 0.27 0.44 0.00 0.86 0.19 0.46 0.37
AEDA 0.02 0.09 0.17 0.95 0.00 0.14 0.23
BT 0.38 0.42 0.70 0.54 0.36 0.59 0.50
Unmask 0.09 0.11 0.69 0.83 0.05 0.25 0.33
AugGPT 0.23 0.19 0.92 0.47 0.24 0.31 0.39
Grammar 0.64 0.62 1.00 0.00 0.13 0.54 0.49
Spelling 0.14 0.17 0.47 0.83 0.49 0.37 0.41
Chain 0.27 0.19 0.98 0.95 0.48 0.67 0.59
Hint 0.56 0.51 0.98 0.86 0.45 0.68 0.67
Taboo 0.26 0.18 0.94 1.00 0.35 0.59 0.55
DoAug 1.00 1.00 0.87 0.98 1.00 1.00 0.98
Table 2: 6 diversity metrics averaged on 12 datasets and the average score, normalized to $[0,1]$.